Define A = 5i + 6j - 7k and B = 2i - 4j +2k.

Find |A x B|.


1
Expert's answer
2021-09-02T10:41:28-0400

The vector product is given by

"{\bf A\times B}=\begin{vmatrix}
 {\bf i} & {\bf j }& {\bf k} \\
 A_x & A_y &A_z\\
B_x & B_y &B_z
\end{vmatrix}"

"{\bf A\times B}=\begin{vmatrix}
 {\bf i} & {\bf j }& {\bf k} \\
 5 & 6 &-7\\
2 & -4 &2
\end{vmatrix}"

"=-16{\bf i}-24{\bf j}-32{\bf k}"

Finally,

"|{\bf A\times B}|=\sqrt{(-16)^2+(-24)^2+(-32)^2}=43"